# XUE: The CO2-rich terrestrial planet-forming region of an externally irradiated Herbig disk

### v1

@@ Abstract
Results. We detect bright CO2 gas with the first simultaneous detection ( 5) of four isotopologues (12CO2, 13CO2, 16O12C18O, 16O12C17O) in a protoplanetary disk. We also detect faint CO emission (2) and the HI Pf line (8). We also place strict upper limits on the water content, finding a total column density 1018 cm-2. The CO2 species trace low gas temperatures (300–370 K) with a range of column densities of 7.4 1017 cm-2 (16O12C17O)–1.3 1020 cm-2 (12CO2) in an equivalent emitting radius of 1.15 au. The emission of 13CO2 is likely affected by line optical depth effects. 16O12C18O and 16O12C17O abundances may be isotopically anomalous compared to the 16O/18O and 16O/17O ratios measured in the interstellar medium and the Solar System.
Conclusions. We propose that the mid-infrared spectrum of XUE 10 is explained by H2O removal either via advection or strong photo-dissociation by stellar UV irradiation, and enhanced local CO2 gas-phase production. Outer disk truncation supports the observed CO2–H2O dichotomy. A CO2 vapor enrichment in 18O and 17O can be explained by means of external UV irradiation and early on (104-5 yr) delivery of isotopically anomalous water ice to the inner disk.
@@
The discovery of more than 5000 exo-planets in our Galaxy, whose number steadily increases1111https://exoplanetarchive.ipac.caltech.edu/, and the growing evidence for several classes of them (e.g. Super-Earths, Hot Jupiters) differing from the Solar System architecture demand a deep understanding of the properties of Proto-Planetary Disks (PPDs) and their connection to the observed exoplanet demography (Winn & Fabrycky, 2015; Mordasini et al., 2012; Dawson & Johnson, 2018). Our knowledge on the formation of planetary systems is mostly based on surveys of PPDs located in nearby ( 400 pc), relatively isolated star-forming regions (e.g., Taurus; Manara et al., 2023), where it is possible to achieve high angular and spectral resolution coupled with high sensitivity to study in detail the gas and the dust components of disks and their physical structure.
@@
Nevertheless, observations (e.g., Stolte et al., 2010; Guarcello et al., 2023) and theory (Fatuzzo & Adams, 2008; Winter et al., 2020) have come to agree that the majority of stars and planetary systems actually form in massive clusters (e.g. Lada & Lada, 2003; Krumholz et al., 2019). In these star-forming regions, the PPDs are exposed to the external FUV radiation of massive young stars of spectral type O and B present in the surroundings (Winter & Haworth, 2022; Allen et al., 2025). The typical FUV flux (measured between 912 and 2400 ) of irradiated protoplanetary disks ranges (Winter & Haworth, 2022), where 1.6 10-3 erg s-1 cm-2 is the average value measured in the Solar neighborhood (Habing, 1968). The external photoevaporation induced by FUV photons is thought to be the dominant dispersal mechanism of PPDs in massive star-forming regions (Scally & Clarke, 2001; Guarcello et al., 2023; Winter et al., 2018). This process drives significant disk mass-loss rates (Adams et al., 2004; Facchini et al., 2016; Haworth & Clarke, 2019; Haworth et al., 2018, 2023), up to about 10-6 ⊙ yr-1 in the Orion Nebula Cluster (ONC) (Johnstone et al., 1998; Henney & O’Dell, 1999; Ballering et al., 2023; Aru et al., 2024), reducing the gas mass and shortening its dissipation timescale, resulting in a rapid outer disk dispersal (Adams et al., 2004; Concha-Ramírez et al., 2019; Winter et al., 2020; Coleman & Haworth, 2022). Depending on the disk structure, consequent steep surface density gradients in the outer disc may result in rapid inward drift of large dust grains, while small dust grains may be depleted by entrainment in the photoevaporative wind (Facchini et al., 2016; Sellek et al., 2020; Qiao et al., 2023; Gárate et al., 2024; Paine et al., 2025). This depletion of gas and dust may alter the assembly of both rocky and gaseous types of planets, although it might also instigate planet formation by streaming instability, enhancing the local dust-to-gas ratio in the outer disk (Throop & Bally, 2005).
@@
There is compelling evidence for outer disk depletion by external photoevaporation in Orion (e.g., Mann et al., 2014; Ansdell et al., 2017; van Terwisga et al., 2019, 2020) and in other closeby massive clusters (Stolte et al., 2010; Roccatagliata et al., 2011; Guarcello et al., 2023; Richert et al., 2018). Theoretical works have illustrated how chemistry is expected to be affected in the outer and inner disk regions (Walsh et al., 2013; Ndugu et al., 2024; Keyte & Haworth, 2025; Gross & Cleeves, 2025). We identify the inner disk as the region extending out to 10 au from the central pre-main sequence star (PMS). Given the typical gas ( 100–6000 K) and dust temperatures ( 200–1500 K) of the inner disk (Kamp & Dullemond, 2004), near- to mid-infrared (IR) observations are needed to probe this region. Ground based near-IR, the Infrared Space Observatory (ISO), the Spitzer Space Telescope, and now JWST spectroscopic surveys of nearby T Tauri disks (see e.g., Salyk et al., 2008; Pontoppidan et al., 2010; Pascucci et al., 2013; Dishoeck et al., 2023; Henning et al., 2024; Arulanantham et al., 2025) have revealed a plethora of ro-vibrational and pure rotational emission lines of several molecular species, including water, carbon monoxide (CO), carbon dioxide (CO2), hydroxide (OH), and molecular hydrogen, which are all connected to the protoplanetary disk physical structure, its evolution, and the mass budget for forming planets. Thanks to its superior sensitivity and spatial resolution compared to Spitzer, JWST makes it possible to extend this kind of spectroscopic studies to disk populations located at distances 1.5 kpc, therefore including a large sample of typical massive star-forming regions. In the ONC, JWST observations towards the low-mass d203-506 proplyd, i.e., cometary shaped photo-evaporating disk, reported the first ever detection of the methyl cation molecule (CH3+; Berné et al., 2023; Zannese et al., 2025), which was later discovered in the non-irradiated T Tauri TW Hya disk (Henning et al., 2024). In the same proplyd, evidence was also found for the reprocessing of water into OH by means of strong UV photo-dissociation at the disk surface (Zannese et al., 2024). However, it is still unclear how the external photoevaporation influences the inner disk from an evolutionary and observational perspective.
@@
The eXtreme UV Environments (XUE) program was proposed in order to characterize the physical and chemical properties of a sample of 12 externally irradiated PPDs in the NGC 6357 massive cluster, taking advantage of pivotal mid-IR JWST Cycle 1 observations obtained with the Medium Resolution Spectrometer (MRS) of the Mid-InfraRed Instrument (MIRI; Ramirez-Tannus et al., 2021). NGC 6357 represents one of the youngest (0.8–1.4 Myr) and closest (1.69 kpc) massive clusters now accessible with JWST (Fang et al., 2012; Getman et al., 2014; Ramírez-Tannus et al., 2020, 2023), hosting more than 20 O stars, and also one of the most massive stars in our Galaxy, Pis24 1 (⋆ 74 ⊙, O4III(f+)+O3.5If*; Massey et al., 2001; Walborn et al., 2002). NGC 6357 is constituted by three coeval subclusters (Pismis 24, G353.1+0.6, and G353.2+0.7) located at the same distance, but where it is possible to probe different ranges of FUV field strenghts, from approximately 103 0 (in G353.2+0.7) to 106 0 (in Pismis 24). Ramírez-Tannus et al. (2023) presented the first results on the XUE 1 solar analogue, which retains similar inner disk properties to nearby non-irradiated T Tauri. Portilla-Revelo et al. (2025) recently validated these findings with a thermo-chemical disk model, and found evidence for a truncated disk with outer gas depletion. Ramírez-Tannus et al. (2025) recently reviewed the general properties of the XUE sources with respect to other samples, while a forthcoming work will present radiation thermo‑chemical simulations tailored to the UV field in NGC 6357 (Hernández et al., in prep.).
@@
Except for XUE 1, the XUE sample is constituted by disks surrounding Intermediate Mass T Tauri (IMTTs) stars. These are characterized by higher stellar masses than T Tauri sources (⋆ 1.5–10 ⊙), with a late spectral type F and G, and will eventually become Herbig A stars (Valegård et al., 2021; Brittain et al., 2023). Herbig stars have earlier spectral type A and B, and represent the prototypical formation sites of giant planets (Johnson et al., 2010; Reffert et al., 2015). This work is dedicated to the spectral analysis of the XUE 10 disk around a early F-type IMTT star, having an effective temperature of about 7000 K, and a stellar mass 3 . This makes XUE 10 a borderline case between a IMTT and an Herbig star, as shown by Valegård et al. (2021). We classify XUE 10 as a disk surrounding an Herbig star. XUE 10 has coordinates R.A. (J2000) = 17h25m55s.274 and Dec. (J2000) = –34∘15’47”.76, and is located in the G353.2+0.7 subcluster of NGC 6357 with an estimated FUV exposure of 5.2 103 G0 (Ramírez-Tannus et al., 2025).
@@
The remainder of the manuscript is organized as follows. The data used for the analysis in this work and the adopted data reduction are presented in Section 2. In Section 3 we present the results on the gas molecular chemistry and the central star. In Section 4 we determine the main physical parameters of gas-phase molecules by means of 0D LTE slab models. In Section 5 we place our findings in context with the available disk samples with similar infrared characterization, and we discuss different scenarios to explain the observed CO2-rich mid-IR spectrum of XUE 10. The conclusions are summarized in Section 6.
@@ 2.1 JWST/MIRI MRS IFU spectroscopy
XUE 10 was observed with the MIRI MRS instrument (Rieke et al., 2015; Wells et al., 2015; Wright et al., 2015) on the 13th of September 2023 as part of the XUE project in Cycle 1 (GO-1759; Ramirez-Tannus et al., 2021). All three MIRI wavelength settings (SHORT, MEDIUM, and LONG) were used for the observations. A four-point dither optimized for a point source was performed in the negative direction. The observations were obtained in the FASTR1 readout mode with a total exposure time of 2700 seconds. No dedicated background observations were taken, as the background nebula in the observations is bright and highly spatially variable.
@@ 2.2 VLT/FORS2 slit spectroscopy
The FORS2 instrument at the VLT, working in the 3300–11000 wavelength range, was used in Long-Slit (LSS) spectroscopic mode to observe XUE 10 on the 20th April 2024 as part of program 113.26CR (P.I.: A. Bik). The observations were obtained with an airmass of 1.06 using a 07 wide slit and the Grism 1028z (GRIS_1028z+29), which observes at 7730–9480 with a resolving power 2560.
@@ 3.1 MIRI MRS spectrum
At the mid-infrared wavelengths covered by the MIRI MRS instrument, the emission shown in Figure 1 comes from the inner disk regions of XUE 10. The spectrum shows a rising dust continuum, indicative of the re-emission at infrared wavelengths of heated dust grains in the upper layers of the disk. We detect emission from gas-phase polycyclic aromatic hydrocarbon (PAHs) particles at 6.2 m and 11.3 m. At the latter wavelength, the PAHs feature is overlapping with an emission band of forsterite (van Boekel et al., 2005). There is also evidence for the PAHs complex at 7.7–8.6 m, but this coincides with the blue shoulder of the silicate band, and therefore does not clearly stand out. We also detect a broad emission feature between 15.5 and 17 m, whose origin remains unclear. We discuss this feature further in Appendix B.
@@
Regarding the molecular gas-phase component, we detect unusually bright CO2 emission. We detect for the first time in a protoplanetary disk spectrum four isotopic species, i.e., isotopologues of CO2, between 12.9 and 17.6 m at high significance ( 5) above the continuum level: 12CO2, 13CO2, 16O12C18O, and 16O12C17O. We do not only detect the fundamental -branch of these isotopes (corresponding to the 000 010 transition of the bending mode), but also hot -branch bandheads (e.g., 011 111) and blended - and -branch lines (see Figure 3). CO2 is a commonly detected species in protoplanetary disks, and is one of the crucial diagnostics for their chemo-physical structure and the processes shaping them (e.g., Öberg et al., 2011; Pontoppidan & Blevins, 2014). The isotopes are naturally less abundant, and therefore fainter and harder to detect. Besides XUE 10, the 13CO2 isotopologue has been observed by now in several nearby T Tauri disks (Grant et al., 2023; Gasman et al., 2025). Vlasblom et al. (2025) were instead the first to report a tentative detection of 16O12C18O in CX Tau (T Tauri), yet finding no evidence for the 16O12C17O isotopologue. Most recently, Salyk et al. (2025) claimed the first firm detection of 16O12C18O, and also a tentative detection of 16O12C17O in MY Lup, another T Tauri disk. At 4.9–6 m, the MIRI instrument covers the upper side of the ro-vibrational -branch ladder ( 25) of CO, the primary gas molecular tracer in protoplanetary disks. In this wavelength range we detect weak CO emission with a 2 significance on the brightest lines (see Figure 5). We do not find instead robust evidence of water emission in XUE 10. The mid-IR ro-vibrational (5–8 m) and rotational transitions ( 10 m) (Meijerink et al., 2009) of water are thought to probe progressively further regions in the inner disk, with hot and lukewarm components ( 400 K) at the shorter wavelengths, and a cold component ( 200–400 K) arising at longer wavelengths (Banzatti et al., 2020, 2023). We only place an upper limit on water lines at 6–8 m, 17.2 m and 23.9 m (see Figure 6). In the 6–8 m window we serendipitously detect (8) the Pfund line of atomic hydrogen. No other hydrogen lines or other species (e.g., OH, C2H2, HCN) are identified in the mid-IR spectrum of XUE 10.

### v2

@@ Abstract
Results. We detected bright CO2 gas with the first simultaneous detection ( 5) of four isotopologues (12CO2, 13CO2, 16O12C18O, 16O12C17O) in a protoplanetary disk. We also detected faint CO emission (2) and the HI Pf line (8). We placed strict upper limits on the water content, finding a total column density of 1018 cm-2. The CO2 species trace low gas temperatures (300–370 K) with a range of column densities of 7.4 1017 cm-2 (16O12C17O)–1.3 1020 cm-2 (12CO2) in an equivalent emitting radius of 1.15 au. The emission of 13CO2 is likely affected by line optical depth effects. The 16O12C18O and 16O12C17O abundances may be isotopically anomalous compared to the 16O/18O and 16O/17O ratios measured in the interstellar medium and the Solar System.
Conclusions. We propose that the mid-infrared spectrum of XUE 10 is explained by H2O removal either via advection or strong photo-dissociation by stellar UV irradiation and enhanced local CO2 gas phase production. Outer disk truncation supports the observed CO2–H2O dichotomy. A CO2 vapor enrichment in 18O and 17O can be explained by means of external UV irradiation and early (104-5 yr) delivery of isotopically anomalous water ice to the inner disk.
@@
The discovery of more than 5000 exo-planets in our Galaxy, a number that steadily increases,1111https://exoplanetarchive.ipac.caltech.edu/ and the growing evidence for several classes of them (e.g., super-Earths, hot Jupiters) differing from the Solar System architecture demand a deep understanding of the properties of protoplanetary disks and their connection to the observed exoplanet demography (Winn & Fabrycky, 2015; Mordasini et al., 2012; Dawson & Johnson, 2018). Our knowledge on the formation of planetary systems is mostly based on surveys of protoplanetary disks located in nearby ( 400 pc), relatively isolated star-forming regions (e.g., Taurus; Manara et al., 2023), where it is possible to achieve the high angular and spectral resolution coupled with high sensitivity needed to study the gas and the dust components of disks in detail as well as their physical structure.
@@
Nevertheless, observations (e.g., Stolte et al., 2010; Guarcello et al., 2023) and theory (Fatuzzo & Adams, 2008; Winter et al., 2020) have come to agree that the majority of stars and planetary systems actually form in massive clusters (e.g., Lada & Lada, 2003; Krumholz et al., 2019). In these star-forming regions, the protoplanetary disks are exposed to the external far-ultraviolet (FUV) radiation of massive young stars of spectral type O and B present in the surroundings (Winter & Haworth, 2022; Allen et al., 2025). The typical FUV flux (measured between 912 and 2400 ) of irradiated protoplanetary disks is in the range of (Winter & Haworth, 2022), where 1.6 10-3 erg s-1 cm-2 is the average value measured in the solar neighborhood (Habing, 1968). The external photoevaporation induced by FUV photons is thought to be the dominant dispersal mechanism of protoplanetary disks in massive star-forming regions (Scally & Clarke, 2001; Guarcello et al., 2023; Winter et al., 2018). This process drives significant disk mass-loss rates (Adams et al., 2004; Facchini et al., 2016; Haworth & Clarke, 2019; Haworth et al., 2018, 2023), up to about 10-6 ⊙ yr-1 in the Orion Nebula Cluster (ONC; Johnstone et al., 1998; Henney & O’Dell, 1999; Ballering et al., 2023; Aru et al., 2024), reducing the gas mass, shortening its dissipation timescale, and resulting in a rapid outer disk dispersal (Adams et al., 2004; Concha-Ramírez et al., 2019; Winter et al., 2020; Coleman & Haworth, 2022). Depending on the disk structure, consequent steep surface density gradients in the outer disc may result in rapid inward drift of large dust grains, while small dust grains may be depleted by entrainment in the photoevaporative wind (Facchini et al., 2016; Sellek et al., 2020; Qiao et al., 2023; Gárate et al., 2024; Paine et al., 2025). Gas and dust depletion may alter the assembly of both rocky and gaseous types of planets, although it might also instigate planet formation by streaming instability, enhancing the local dust-to-gas ratio in the outer disk (Throop & Bally, 2005).
@@
There is compelling evidence for outer disk depletion by external photoevaporation in Orion (e.g., Mann et al., 2014; Ansdell et al., 2017; van Terwisga et al., 2019, 2020) and in other close by massive clusters (Stolte et al., 2010; Roccatagliata et al., 2011; Guarcello et al., 2023; Richert et al., 2018). Theoretical works have illustrated how chemistry is expected to be affected in the outer and inner disk regions (Walsh et al., 2013; Ndugu et al., 2024; Keyte & Haworth, 2025; Gross & Cleeves, 2025). We identify the inner disk as the region extending out to 10 au from the central pre-main-sequence star (PMS). Given the typical gas ( 100–6000 K) and dust temperatures ( 200–1500 K) of the inner disk (Kamp & Dullemond, 2004), near- to mid-infrared (IR) observations are needed to probe this region. Ground-based near-IR, the Infrared Space Observatory, the Spitzer Space Telescope, and now JWST spectroscopic surveys of nearby T Tauri disks (see, e.g., Salyk et al., 2008; Pontoppidan et al., 2010; Pascucci et al., 2013; Dishoeck et al., 2023; Henning et al., 2024; Arulanantham et al., 2025) have revealed a plethora of ro-vibrational and pure rotational emission lines of several molecular species, including water, carbon monoxide (CO), carbon dioxide (CO2), hydroxide (OH), and molecular hydrogen, which are all connected to the physical structure of a protoplanetary disk, its evolution, and the mass budget for forming planets. Thanks to its superior sensitivity and spatial resolution compared to Spitzer, JWST makes it possible to extend this kind of spectroscopic study to disk populations located at distances 1.5 kpc, therefore including a large sample of typical massive star-forming regions. In the ONC, JWST observations toward the low-mass d203-506 proplyd, i.e., a cometary-shaped photo-evaporating disk, reported the first ever detection of the methyl cation molecule (CH3+; Berné et al., 2023; Zannese et al., 2025), which was later discovered in the non-irradiated T Tauri TW Hya disk (Henning et al., 2024). In the same proplyd, evidence was also found for the reprocessing of water into OH by means of strong UV photo-dissociation at the disk surface (Zannese et al., 2024). However, it is still unclear how the external photoevaporation influences the inner disk from an evolutionary and observational perspective.
@@
The eXtreme UV Environments (XUE) program was proposed in order to characterize the physical and chemical properties of a sample of 12 externally irradiated protoplanetary disks in the NGC 6357 massive cluster, taking advantage of pivotal mid-IR JWST Cycle 1 observations obtained with the Medium Resolution Spectrometer (MRS) of the Mid-InfraRed Instrument (MIRI; Ramirez-Tannus et al., 2021). NGC 6357 represents one of the youngest (0.8–1.4 Myr) and closest (1.69 kpc) massive clusters now accessible with JWST (Fang et al., 2012; Getman et al., 2014; Ramírez-Tannus et al., 2020, 2023), hosting more than 20 O stars and one of the most massive stars in our Galaxy: Pis24 1 (⋆ 74 ⊙, O4III(f+)+O3.5If*; Massey et al., 2001; Walborn et al., 2002). NGC 6357 comprises three coeval subclusters (Pismis 24, G353.1+0.6, and G353.2+0.7) located at the same distance, but it is possible to probe different ranges of FUV field strengths, from approximately 103 0 (in G353.2+0.7) to 106 0 (in Pismis 24). Ramírez-Tannus et al. (2023) presented the first results on the XUE 1 solar analog, which retains similar inner disk properties to nearby non-irradiated T Tauri. Portilla-Revelo et al. (2025) recently validated these findings with a thermo-chemical disk model and found evidence for a truncated disk with outer gas depletion. Ramírez-Tannus et al. (2025) recently reviewed the general properties of the XUE sources with respect to other samples, while a forthcoming work will present radiation thermo-chemical simulations tailored to the UV field in NGC 6357 (Hernández et al., in prep.).
@@
Except for XUE 1, the XUE sample is constituted by disks surrounding intermediate mass T Tauri (IMTTs) stars. These are characterized by higher stellar masses than T Tauri sources (⋆ 1.5–10 ⊙), with a late spectral type F and G, and they will eventually become Herbig A stars (Valegård et al., 2021; Brittain et al., 2023). Herbig stars have an earlier spectral type, A and B, and they represent the prototypical formation sites of giant planets (Johnson et al., 2010; Reffert et al., 2015). This work is dedicated to the spectral analysis of the XUE 10 disk around a early F-type IMTT star with an effective temperature of about 7000 K and a stellar mass of 3 . This makes XUE 10 a borderline case between an IMTT and an Herbig star, as shown by Valegård et al. (2021). We classify XUE 10 as a disk surrounding an Herbig star. The source has coordinates R.A. (J2000) = 17h25m55s.274 and Dec. (J2000) = –34∘15’47”.76 and is located in the G353.2+0.7 subcluster of NGC 6357 with an estimated FUV exposure of 5.2 103 G0 (Ramírez-Tannus et al., 2025).
@@
The remainder of the manuscript is organized as follows. The data used for the analysis in this work and the adopted data reduction are presented in Section 2. In Section 3 we present the results on the gas molecular chemistry and the central star. In Section 4 we determine the main physical parameters of gas phase molecules by means of 0D local thermodynamic equilibrium (LTE) slab models. In Section 5 we place our findings in context with the available disk samples with similar infrared characterization, and we discuss different scenarios to explain the observed CO2-rich mid-IR spectrum of XUE 10. The conclusions are summarized in Section 6.
@@ 2.1 JWST/MIRI MRS IFU spectroscopy
XUE 10 was observed with the MIRI MRS instrument (Rieke et al., 2015; Wells et al., 2015; Wright et al., 2015) on the 13 September 2023 as part of the XUE project in Cycle 1 (GO-1759; Ramirez-Tannus et al., 2021). All three MIRI wavelength settings (SHORT, MEDIUM, and LONG) were used for the observations. A four-point dither optimized for a point source was performed in the negative direction. The observations were obtained in the FASTR1 readout mode with a total exposure time of 2700 seconds. No dedicated background observations were taken, as the background nebula in the observations is bright and highly spatially variable.
@@ 2.2 VLT/FORS2 slit spectroscopy
The FORS2 instrument at the VLT, working in the 3300–11000 wavelength range, was used in Long-Slit (LSS) spectroscopic mode to observe XUE 10 on the 20 April 2024 as part of program 113.26CR (P.I.: A. Bik). The observations were obtained with an airmass of 1.06 using a 07 wide slit and the Grism 1028z (GRIS_1028z+29), which observes at 7730–9480 with a resolving power 2560.
@@ 3.1 MIRI MRS spectrum
At the mid-infrared wavelengths covered by the MIRI MRS instrument, the emission shown in Figure 1 comes from the inner disk regions of XUE 10. The spectrum shows a rising dust continuum, indicative of the re-emission at infrared wavelengths of heated dust grains in the upper layers of the disk. We detect emission from gas phase polycyclic aromatic hydrocarbon (PAH) particles at 6.2 m and 11.3 m. At the latter wavelength, the PAH feature overlaps with an emission band of forsterite (van Boekel et al., 2005). There is also evidence for the PAH complex at 7.7–8.6 m, but this coincides with the blue shoulder of the silicate band and therefore does not clearly stand out. We also detect a broad emission feature between 15.5 and 17 m, whose origin remains unclear. We discuss this feature further in Appendix B.
@@
Regarding the molecular gas phase component, we detect unusually bright CO2 emission. We detect for the first time in a protoplanetary disk spectrum four isotopic species, i.e., isotopologues of CO2, between 12.9 and 17.6 m at high significance ( 5) above the continuum level: 12CO2, 13CO2, 16O12C18O, and 16O12C17O. We do not only detect the fundamental -branch of these isotopes (corresponding to the 000 010 transition of the bending mode), but also hot -branch bandheads (e.g., 011 111) and blended - and -branch lines (see Figure 3). CO2 is a commonly detected species in protoplanetary disks, and is one of the crucial diagnostics for their chemo-physical structure and the processes shaping them (e.g., Öberg et al., 2011; Pontoppidan & Blevins, 2014). The isotopes are naturally less abundant, and therefore fainter and harder to detect. Besides XUE 10, the 13CO2 isotopologue has been observed by now in several nearby T Tauri disks (Grant et al., 2023; Gasman et al., 2025). Vlasblom et al. (2025) were instead the first to report a tentative detection of 16O12C18O in CX Tau (T Tauri), yet finding no evidence of the 16O12C17O isotopologue. Most recently, Salyk et al. (2025) claimed the first firm detection of 16O12C18O, and also a tentative detection of 16O12C17O in MY Lup, another T Tauri disk. At 4.9–6 m, the MIRI instrument covers the upper side of the ro-vibrational -branch ladder ( 25) of CO, the primary gas molecular tracer in protoplanetary disks. In this wavelength range we detect weak CO emission with a 2 significance on the brightest lines (see Figure 5). We do not find instead robust evidence of water emission in XUE 10. The mid-IR ro-vibrational (5–8 m) and rotational transitions ( 10 m) (Meijerink et al., 2009) of water are thought to probe progressively further regions in the inner disk, with hot and lukewarm components ( 400 K) at the shorter wavelengths, and a cold component ( 200–400 K) arising at longer wavelengths (Banzatti et al., 2020, 2023). We only place an upper limit on water lines at 6–8 m, 17.2 m and 23.9 m (see Figure 6). In the 6–8 m window we serendipitously detect (8) the Pfund line of atomic hydrogen. No other hydrogen lines or other species (e.g., OH, C2H2, HCN) are identified in the mid-IR spectrum of XUE 10.